A student pulls a 10 N block along a rough horizontal surface. The horizontal and the vertical components of the force that he applies are 4 N and 3 N, respectively.
What is the normal force?
1
Expert's answer
2019-10-28T10:57:08-0400

There is the equilibrium in the vertical direction:


"N+F_y-W=0"

The normal force:


"N=W-F_y=10-3=7 \ N"